# Masking Augmentation for Supervised Learning

### v2

@@ 3.3 Expand to drop regularizations
We design MaskSub for masking augmentation. Due to its simplicity, it can be expanded to drop-based regularizations [44, 17, 27]. In this section, we introduce two variants of MaskSub: DropSub for dropout [44] and PathSub for drop-path [27]. Since the drop-based regularizations easily adjust their strength by controlling drop probability, MaskSub enables the model to learn strong drop regularization without degrading performance at standard loss, similar to masking augmentation. The performance of MaskSub variants is shown in Section 4.7 and A.4.
DropSub. Dropout [44] is a fundamental drop regularization. Dropout drops random elements of network features with a fixed probability. Since dropout is not related to feature structure, every element of the feature has independent drop probability $p_{drop}$. DropSub is simply implemented by changing $r_{mask}$ to the dropout probability $p_{drop}$. Thus, the sub-model uses strong dropout, while the main-model follows a standard training recipe. Due to stability issues, dropout is not preferred in recent training recipes [45, 47]. However, DropSub enables strong dropout in ViT training and achieves performance improvement.
PathSub. Drop-path [27, 16] randomly drops a total feature of the network block with a probability $p_{path}$. PathSub is also implemented by changing $r_{mask}$ to the drop-path probability $p_{path}$. Drop-path widely used in training recipes [45, 47, 54] to adjust the regularization strength [47]. Thus, different from previous cases, the main-model uses the drop-path following the training recipe, and the sub-model uses a higher drop probability than the main-model.
4 Experiments
We validate the effectiveness of Masked Sub-model (MaskSub) by applying it to diverse training scenarios. We claim MaskSub is an easy plug-in solution for various training recipes. Thus, we strictly follow the original training recipe, including optimizer parameters, learning rate and weight-decay, and regularization parameters. The only difference between baseline and MaskSub is the masking augmentation for the sub-model. We simply set the masking ratio of MaskSub to 50% across all experiments. In short, MaskSub does not have a hyper-parameter that varies depending on training scenarios.
4.1 Training from scratch
The training recipe in ViTs is a key factor enabling ViT to surpass CNN; thus, the ViT training recipe is a significant and active research topic. We use a state-of-the-art ViT recipe, DeiT-III [47], as a baseline. Integrating additional techniques into the DeiT-III is a significant challenge, and improvements made over DeiT-III can be considered a novel state-of-the-art in ViT training.
@@ 4.1 Training from scratch
Table 1: Training from scratch with ViT using DeiT-III. MaskSub (50%) is applied to the ViT training recipe [47] on ImageNet-1k. Note that the training settings are identical to the original ones.
Network Method Epochs Top-1 acc.
ViT-S DeiT [45] 300 79.8
MAE [${}^{\dagger}$ 1600 81.4
DeiT-III [47] 800 81.4
DeiT-III [47] + CoSub [48] 800 81.5
DeiT-III [47] + MaskSub 800 81.7
ViT-B DeiT [45] 300 81.8
MAE [21] 1600 83.6
DeiT-III [47] 800 83.8
SupMAE [34] 400 83.6
DeiT-III [47] + CoSub [48] 800 84.2
DeiT-III [47] + MaskSub 400 84.1
DeiT-III [47] + MaskSub 800 84.2
ViT-L DeiT-III [47] 800 84.9
DeiT-III [47] + CoSub [48] 800 85.3
DeiT-III [47] + MaskSub 400 85.2
DeiT-III [47] + MaskSub 800 85.3
ViT-H DeiT-III [47] 800 85.2
DeiT-III [47] + CoSub [48] 800 85.7
DeiT-III [47] + MaskSub 400 85.7
Table 2: Comparison with pre-training methods. We compared DeiT-III [47] + MaskSub with various pre-training methods. MaskSub shows remarkable performance in all sizes.
Various sizes of ViTs are trained with MaskSub (50%) on 400 and 800 epochs training. The results are shown in Table 1. MaskSub significantly improves performance in all settings. In 400 epochs training, MaskSub improves DeiT-III with substantial margins, which even outperforms 800 epochs DeiT-III except for ViT-S/16. MaskSub also demonstrates superior performance in 800 epochs. The impact of MaskSub is impressively sustained even for larger models like ViT-L/16 and ViT-H/16. It is worth noting that ViT-H + MaskSub (400 epochs) outperforms ViT-H/16 (800 epochs) with a significant +0.5pp gain even with half training length. Thus, MaskSub is an effective way to improve ViT training.
Table 2 shows performances of MaskSub compared to other pre-training methods. With a reasonable performance gap, MaskSub outperforms MAE [21] in ViT-S and ViT-B. Compared to SupMAE [34], MaskSub achieves better performance with the same training epoch. CoSub [48] has comparable performance with MaskSub but, as shown in Section 4.4, MaskSub requires less computation costs than CoSub. Thus, we argue that MaskSub outperforms CoSub.
@@ 4.2 Finetuning
Table 3: ImageNet-1k finetuning. We report finetuned performance of MAE [21], BEiT v2 [38] and CLIP finetuning [14] with MaskSub (50%). Official pretrained weights are used.
@@
Following the emergence of self-supervised learning [21] and visual-language model [41], the significance of finetuning has notably increased. Generally, self-supervised learning, such as MAE [21] and BEiT [1, 38], does not use supervised labels at pretraining, which makes MaskSub inapplicable for pretraining. However, most methods utilize supervised finetuning steps after pretraining to demonstrate their performance. Thus, we apply our MaskSub (50%) to the finetuning stage to verify the effect of MaskSub on finetuning. Note that we strictly follow original recipes and apply MaskSub (50%) based on it. All finetuning is conducted using officially released pretrained weights.
We utilize three finetuning recipes: MAE [21], BEiT v2 [38], and Finetune CLIP [14]. MAE [21] is a representative method of masked image models (MIM). Since our random masking is motivated by MAE, MaskSub is seamlessly integrated into MAE finetuning process. BEiT v2 [38] utilizes pretrained CLIP for MIM and achieves superior performance compared to MAE. Following the masking strategy of BEiT v2 using mask-token, we adjust MaskSub to masking using mask-token from the pretrained weight instead of MAE-style masking. Finetune CLIP [14] is a finetuning recipe for CLIP [41] pretrained weights. MaskSub is applied to finetuning CLIP without change.
Table 3 shows the finetuning results. MaskSub improves the performance of all finetune recipes, including large-scale ViT models. This is notable as it shows substantial improvement with a short finetuning phase of fewer than 100 epochs compared to the pretraining period of 1600 epochs. In MAE finetuning, MaskSub improves 0.2 - 0.3pp in all model sizes. MaskSub is also effective on BEiT v2, which utilizes RPE [36, 1] and masking strategy with mask-token. Even in CLIP finetuning, MaskSub achieves substantial improvement. In finetuning CLIP, we report performance at the last epoch rather than selecting the best performance in early epochs. The best performance of finetuning CLIP with MaskSub is the same as the baseline. Table 4 demonstrates the impacts of MaskSub compared to cutting-edge CLIP-based training recipes. It shows that MaskSub improves the performance of state-of-the-art training recipes.
@@ 4.3 Hierarchical architecture
Table 6: Comparison with self-distillation methods. Based on ResNet performance, we compare MaskSub with various self-distillation methods.
We extend experiments to architectures with hierarchical spatial dimensions: ResNet [19] and Swin Transformer [36]. Unlike ViT maintains spatial token length for all layers, those networks change the spatial size of features in the middle of layers, requiring a change in masking strategy. We apply MaskSub (50%) to ResNet and Swin Transformer. We simply fill out masked regions with zero pixels for ResNets and replace masked regions with mask-tokens for Swin Transformer. It maintains the spatial structure and enables spatial size reduction of hierarchical architecture. Following a literature [55], we use random masking with patch-size $32\times 32$. Note that the computation reduction using MAE-style masking is not applicable for this case. Thus, MaskSub costs double the training budget. For ResNet, we use a high-performance training recipe [54] with 300 epochs. The recipe of original paper [36] is used for the Swin Transformer training. We strictly follow the training recipe and apply MaskSub without recipe tuning.
Results are shown in Table 5. MaskSub achieves impressive performance gains even in ResNet and Swin Transformer. ResNet and Swin are substantially different architectures from ViT. Thus, the result implies that the effectiveness of MaskSub is not limited to ViT architectures and is applicable to hierarchical architectures. Using ResNet results, we compare MaskSub with various self-distillation methods. As shown in Table 6, most of the self-distillations report performance based on weak training recipes. Thus, they are incompatible with cutting-edge training recipes such as RSB [54], which limits the general applicability of self-distillation. Otherwise, MaskSub can easily be plugged into strong training recipes and contribute to achieving state-of-the-art performance.
4.4 Training budget
@@
Table 7: Comparison in the same training budget. All training has been conducted with NVIDIA V100 8 GPUs. GPU days refer to the number of days required for training when using a single V100 GPU.
We have shown that MaskSub effectively improves the performance of various architectures. However, MaskSub requires additional computation costs for the sub-model, which increases training costs. Thus, we analyze MaskSub regarding its training costs to determine if MaskSub could be an effective solution within a limited training budget. We compare MaskSub with training recipes with increased epochs to align with the training budget. The training budget is quantified regarding required GPU days when only a single NVIDIA V100 GPU is used for training. Table 7 shows the results. In ViT training, MaskSub outperforms baseline with ×$\times 1.5$ epochs setting. Thus, MaskSub is superior to the long epoch training to spend computation costs for training ViT. For ResNet, we compare 300 epochs MaskSub with 600 epochs training recipe RSB [54] A1. MaskSub outperforms 600 epochs training recipes in ResNet101 and ResNet152. Consequently, the results show that MaskSub is an effective way to improve training, even considering computation costs for the sub-model.
@@
Table 8: Comparison in ImageNet-1k training. The table shows performance and GPU costs of augmenting methods for ViT-B 400 epoch training.
Method Accuracy GPU days
Baseline [47] 83.6 6.0
DataAugment [25] 83.1 (-0.5) 11.8 (+97%)
GradAug [60] 84.0 (+0.4) 23.4 (+290%)
CoSub [48] 83.8 (+0.2) 11.0 (+83%)
MaskSub 83.9 (+0.3) 8.6 (+43%)
Table 9: Comparison in MAE finetuning. The table displays performances and GPU costs of augmenting methods when applied to MAE ViT-B fine-tuning.
Model
Pretraining
+ MaskSub
Finetuning
+ MaskSub
CIFAR-10 [31] CIFAR-100 [31] Flowers [37] Cars [30] iNat-18 [49] iNat-19 [49]
ViT-S/16 - - 98.8 90.0 94.5 80.9 70.1 76.7
- 98.9 (+0.1) 90.6 (+0.6) 95.2 (+0.7) 81.2 (+0.3) 70.8 (+0.7) 77.0 (+0.3)
98.8 (+0.0) 89.9 (-0.1) 98.3 (+3.8) 92.2 (+1.3) 71.2 (+1.1) 77.1 (+0.4)
ViT-B/16 - - 99.1 91.7 97.5 90.0 73.2 78.5
- 99.2 (+0.1) 91.9 (+0.2) 97.7 (+0.2) 90.2 (+0.2) 73.6 (+0.4) 78.8 (+0.3)
98.8 (-0.3) 89.6 (-2.1) 98.7 (+1.2) 92.8 (+2.8) 73.9 (+0.7) 79.1 (+0.6)
Table 10: Transfer learning to small-scale datasets. Table shows transfer learning performance with/without MaskSub. We measure the performance when MaskSub is applied to pretraining and finetuning. The standard deviations over three runs are reported in Table A.9.
Table 11: GLUE [50] benchmark with BERT [12]. We apply MaskSub on GLUE benchmark to validate the effect of MaskSub on language model finetuning.
Model +MaskSub MNLI QQP QNLI SST-2 CoLA STS-B MRPC RTE Avg. BERT [12] base - 84.1 87.5 91.0 91.6 54.7 87.0 88.5 62.8 80.9 84.5 87.7 91.3 91.9 58.3 86.8 89.2 63.2 81.6 BERT [12] large - 86.8 88.2 92.3 93.8 63.3 89.3 92.0 69.7 84.4 87.1 89.0 92.7 94.0 65.2 88.6 91.5 69.3 84.7
Training method ImageNet-1k Zero-shot acc. CLIP [41] 33.5 CLIP [41] + Masking 29.8 (-3.7) CLIP [41] + MaskSub 37.6 (+4.1)
Table 11: GLUE [50] benchmark with BERT [12]. We apply MaskSub on GLUE benchmark to validate the effect of MaskSub on language model finetuning.
Table 12: MaskSub on CLIP pretraining with CLIP ViT-B/32 models.
@@
We compare MaskSub with other methods for an additional augmentation: DataAugment [25], GradAug [60], and CoSub [48]. DataAugment [25] uses doubled data augmentations for the same image which is similar to contrastive learning [7, 4]. GradAug [2020] utilizes a network pruning [61] to build sub-network. CoSub introduces a sub-network based on drop-path [27] and uses the sub-network as mutual learning [65]. Table 8 shows the 400 epoch training from scratch result, and Table 9 represents MAE fine-tuning result. Note that GradAug in Table 8 is 200 epochs training result to adjust computation cost similar to other methods. All augmentation methods require additional computation costs. In particular, GradAug spends almost 300% of additional training costs compared to original training. On the other hand, our MaskSub only requires a small amount of additional costs (below 50%), which is a remarkable advantage in training. With the smallest computation, our MaskSub achieves substantial performance improvements. MaskSub performs superior to CoSub in all cases and performs similarly to GradAug in MAE fine-tuning with 1/6 training costs.

### v3

@@ 3.3 Expand to drop regularizations
We design MaskSub for masking augmentation. Due to its simplicity, it can be expanded to drop-based regularizations [47, 18, 28]. In this section, we introduce two variants of MaskSub: DropSub for dropout [47] and PathSub for drop-path [28]. Since the drop-based regularizations easily adjust their strength by controlling drop probability, MaskSub enables the model to learn dropped features without degrading performance at a standard loss, similar to masking augmentation. The performance of MaskSub variants is shown in Section 4.8. Note that detailed experiments with loss convergence for various drop rates are reported in Table A.5 in the Appendix.
DropSub. Dropout [47] is a fundamental drop regularization. Dropout drops random elements of network features with a fixed probability. Since dropout is unrelated to feature structure, every feature element has independent drop probability $p_{drop}$. DropSub is simply implemented by changing $r_{mask}$ to the dropout probability $p_{drop}$. Thus, the sub-branch uses strong dropout, while the main-branch follows a standard training recipe. Due to stability issues, dropout is not preferred in recent training recipes [48, 50]. However, DropSub enables strong dropout in ViT training and achieves performance improvement.
PathSub. Drop-path [28, 17] randomly drops a total feature of the network block with a probability $p_{path}$. PathSub is also implemented by changing $r_{mask}$ to the drop-path probability $p_{path}$. Drop-path widely used in training recipes [48, 50, 57] to adjust the regularization strength [50]. Thus, unlike previous cases, the main-branch uses the drop-path following the training recipe, and the sub-branch uses a higher drop probability than the main-branch.
@@
Table 1: Training from scratch with ViT using the DeiT-III. MaskSub (50%) is applied to the ViT training [50] on ImageNet-1k. Note that the training settings are identical to the original ones.
Network Method Epochs Top-1 acc. Cost
ViT-S DeiT [48] 300 79.8 -
MAE [22] 1600 81.4 -
DeiT-III [50] 800 81.4 ×$\times 1.0$
CoSub [51] 800 81.5 ×$\times 2.0$
MaskSub 400 81.1 ×$\times 0.75$
MaskSub 800 81.7 ×$\times 1.5$
ViT-B DeiT [48] 300 81.8 -
MAE [22] 1600 83.6 -
SupMAE [36] 400 83.6 -
DeiT-III [50] 800 83.8 ×$\times 1.0$
CoSub [51] 800 84.2 ×$\times 2.0$
MaskSub 400 84.1 ×$\times 0.75$
MaskSub 800 84.2 ×$\times 1.5$
ViT-L DeiT-III [50] 800 84.9 ×$\times 1.0$
CoSub [51] 800 85.3 ×$\times 2.0$
MaskSub 400 85.2 ×$\times 0.75$
MaskSub 800 85.3 ×$\times 1.5$
ViT-H DeiT-III [50] 800 85.2 ×$\times 1.0$
CoSub [51] 800 85.7 ×$\times 2.0$
MaskSub 400 85.7 ×$\times 0.75$
Table 2: Pre-training methods comparison. We compare DeiT-III [50] + MaskSub with various pre-training methods. MaskSub shows remarkable performances compared to its training cost.
4 Experiments
We validate the effectiveness of our Masked Sub-branch (MaskSub) by applying it to diverse training scenarios. We claim MaskSub is an easy plug-in solution for various training recipes. Thus, we strictly follow the original training recipe, including optimizer parameters, learning rate and weight-decay, and regularization parameters. The only difference between baseline and MaskSub is the masking augmentation for the sub-branch. We simply set the masking ratio of MaskSub to 50% across all experiments. In short, MaskSub does not have a hyper-parameter that varies depending on training scenarios.
4.1 Training from scratch (pretraining)
The training recipe in ViTs is a key factor enabling ViT to surpass CNN; thus, the ViT training recipe is an important and active research topic. We use a state-of-the-art ViT training recipe, DeiT-III [50], as our baseline. Enhancing DeiT-III by integrating additional techniques is challenging, so we believe improvements made over DeiT-III would represent a new state-of-the-art in ViT training.
ViTs are trained with MaskSub (50%) on 400 and 800 epochs training. The results are shown in Table 1. MaskSub improves performance across all settings. For 400-epochs training, MaskSub improves DeiT-III with substantial margins, which even outperforms 800-epochs trained DeiT-III except for ViT-S/16. MaskSub also demonstrates superior performance when training for 800 epochs. The impact of MaskSub is impressively consistent with larger models like ViT-L/16 and ViT-H/16. It is worth noting that ViT-H + MaskSub (400 epochs) outperforms ViT-H/16 (800 epochs) with +0.5pp gain, even with half the training epochs. Thus, MaskSub is an effective way to improve ViT training.
Table 2 shows the performance and computation cost of MaskSub compared to other pretrainings. In ViT-S and ViT-B, MaskSub outperforms MAE [22] with a reasonable performance gap. Compared to SupMAE [36], MaskSup outperforms under the same epochs. CoSub [51] has comparable performance with MaskSub; however, MaskSub requires less computation costs than CoSub. Thus, we argue that MaskSub outperforms CoSub. More comparisons with CoSub are included in Section 4.5.
@@ 4.2 Finetuning
Table 3: ImageNet-1k finetuning. We report finetuning performance of MAE [22], BEiT v2 [41] and CLIP finetuning [15] with MaskSub (50%). Official weights are used.
Following the emergence of self-supervised learning [22] and visual-language modeling [44], the significance of finetuning has notably increased. Generally, self-supervised learning, such as MAE [22] and BEiT [1, 41], does not use supervised labels at pretraining, which makes MaskSub inapplicable for pretraining. However, a standard is to evaluate the model’s capability using supervised finetuning after pretraining. Thus, we apply our MaskSub (50%) to the finetuning stage to verify the effect of MaskSub on finetuning. Note that we strictly follow the original recipes mentioned below and apply MaskSub (50%) based on it. All finetuning is conducted using officially released pretrained weights.
We utilize three finetuning recipes: MAE [22], BEiT v2 [41], and Finetune CLIP [15]. MAE [22] is a representative method of masked image models (MIM). Since our random masking is motivated by MAE, MaskSub is seamlessly integrated into the MAE finetuning process. BEiT v2 [41] utilizes the pretrained CLIP for MIM and achieves superior performance compared to MAE. Following the masking strategy of BEiT v2 using mask-token, we adjust MaskSub to masking using mask-token from the pretrained weight instead of MAE-style masking. Finetune CLIP [15] is a finetuning recipe for CLIP [44] pretrained weights. MaskSub is applied to finetuning CLIP without change.
@@
Table 3 shows the finetuning results. MaskSub improves the performance of all finetune practices, including large-scale ViT models. This is notable as it shows substantial improvement with a short finetuning phase of fewer than 100 epochs compared to the pretraining period of 1600 epochs. In MAE finetuning, MaskSub improves 0.2 - 0.3pp in all model sizes. MaskSub is also effective on BEiT v2, which utilizes Relative Position Encoding (RPE) [38, 1] and block-masking strategy with mask-tokens. CLIP finetuning also displays that MaskSub achieves substantial improvements. In finetuning CLIP, we report performance at the last epoch rather than selecting the best performance in early epochs. The best performance of finetuning CLIP with MaskSub is the same as the baseline. Table 4 demonstrates the impacts of MaskSub compared to cutting-edge CLIP-based training recipes. It shows that MaskSub improves the performance of state-of-the-art training recipes.
@@ 4.3 Hierarchical architecture
Table 6: Comparison with self-distillation methods. Based on ResNet, we compare MaskSub with self-distillation methods.
We extend experiments to architectures with hierarchical spatial dimensions: ResNet [20] and Swin Transformer [38]. Unlike ViT, which maintains spatial token length for all layers, those networks change the spatial size of features in the middle of layers, requiring a change in masking strategy. We apply MaskSub (50%) to ResNet and Swin Transformer. We simply fill out masked regions with zero pixels for ResNets and replace masked regions with mask-tokens for Swin Transformer. It maintains the spatial structure and enables spatial size reduction of hierarchical architecture. Following the literature [58], we use random masking with the patch size of $32\times 32$. Note that the computation reduction in MAE-style masking does not apply here; therefore, MaskSub costs double the training budget. For ResNet, we use an effective training recipe [57] with 300 epochs. The recipe in the original paper [38] is used for the Swin Transformer training. We strictly follow the training recipes and apply MaskSub without tuning them.
Results are shown in Table 5. MaskSub achieves impressive performance gains with ResNet and Swin Transformer as well. ResNet and Swin are substantially different architectures from ViT. Thus, the result implies that the effectiveness of MaskSub is not limited to ViT architectures and is applicable to hierarchical architectures.
Model
Pretraining
+ MaskSub
Finetuning
+ MaskSub
CIFAR100
[33]
CIFAR100
[33]
Flowers
[40]
Cars
[32]
iNat-18
[52]
iNat-19
[52]
ViT-S/16 - - 98.8 90.0 94.5 80.9 70.1 76.7
- 98.9 90.6 95.2 81.2 70.8 77.0
98.8 89.9 98.3 92.2 71.2 77.1
ViT-B/16 - - 99.1 91.7 97.5 90.0 73.2 78.5
- 99.2 91.9 97.7 90.2 73.6 78.8
98.8 89.6 98.7 92.8 73.9 79.1
Table 7: Transfer learning. Table shows transfer learning performance with/without MaskSub. We measure the performance when MaskSub is applied to pretraining and finetuning. The standard deviations over three runs are reported in Appendix.
@@
Table 8: Comparison in the same training budget. Training has been conducted with NVIDIA V100 8 GPUs. GPU days refer to the number of days required for training when using a V100 GPU.
@@
Table 9: ImageNet-1k Comparison. The table shows performance and computational costs for ViT-B’s 400 epoch training.
Single-scale mIoU Multi-scale mIoU
DeiT-III + MaskSub DeiT-III + MaskSub
ViT-B 48.8 49.4 (+0.6) 49.7 50.2 (+0.5)
ViT-L 51.7 52.2 (+0.5) 52.3 52.7 (+0.4)
Table 10: Semantic segmentation on ADE-20k. UpperNet for ViT backbone is trained with the BEiTv2 segmentation recipe.
4.4 Self-distillation
We compare MaskSub with self-distillation methods. As shown in Table 6, most self-distillations report their performance based on weak and old recipes. Thus, they are less effective with cutting-edge recipes (RSB [57]) or architectures (ViT). Otherwise, MaskSub can be plugged into strong training recipes and achieves state-of-the-art with self-distillation loss. Thus, MaskSub has contributed to practical self-distillation with its broad applicability.
4.5 Training budget
We have shown that MaskSub effectively improves the performance of various architectures. However, MaskSub requires additional computation costs for the sub-branch, which increases training costs. Thus, we analyze MaskSub regarding its training costs to determine if MaskSub could be an effective solution within a limited training budget. We compare MaskSub with training recipes with increased epochs to align with the training budget. The training budget is quantified regarding required GPU days when only a single NVIDIA V100 GPU is used for training. Table 8 shows the results. In ViT training, MaskSub outperforms baseline with ×$\times 1.5$ epochs setting. Thus, MaskSub is superior to the long epoch training to spend computation costs for training ViT. For ResNet, we compare 300 epochs MaskSub with 600 epochs training recipe RSB [57] A1. MaskSub outperforms 600 epochs training recipes in ResNet101 and ResNet152. Consequently, the results show that MaskSub is an effective way to improve training, even considering computation costs for the sub-branch.
@@ 4.5 Training budget
We compare MaskSub with other training methods: DataAugment [26], GradAug [64], and CoSub [51]. DataAugment [26] uses doubled data augmentations for the same image, which is similar to contrastive learning [7, 4]. GradAug [64] utilizes a network pruning [65] to build sub-network. CoSub introduces a sub-network based on drop-path [28] and uses the sub-network as mutual learning [69]. Table 9 shows the 400-epochs training from scratch result. Note that GradAug in Table 9 is a 200-epochs training result to adjust computation cost similar to other methods. All augmentation methods require additional computation costs. In particular, GradAug spends almost 300% of additional training costs compared to original training. On the other hand, our MaskSub only requires a small amount of extra costs (below 50%), which is a remarkable advantage in training. With the smallest computation, our MaskSub achieves substantial performance improvements. MaskSub performs superior to CoSub in all cases.
Model +MaskSub MNLI QQP QNLI SST-2 CoLA STS-B MRPC RTE Average
BERT [13] base - 84.1 87.5 91.0 91.6 54.7 87.0 88.5 62.8 80.9
84.5 87.7 91.3 91.9 58.3 86.8 89.2 63.2 81.6
BERT [13] large - 86.8 88.2 92.3 93.8 63.3 89.3 92.0 69.7 84.4
87.1 89.0 92.7 94.0 65.2 88.6 91.5 69.3 84.7
Table 11: GLUE [53] benchmark with BERT [13]. We apply MaskSub on GLUE benchmark to validate the effect of MaskSub on language model fine-tuning. MaskSub effectively improves BERT finetuning performance.
Training method
ImageNet-1k
Zero-shot acc.
CLIP [44] 33.5
CLIP [44] + Masking 29.8 (-3.7)
CLIP [44] + MaskSub 37.6 (+4.1)
Table 12: MaskSub on CLIP pretraining with ViT-B/32. We apply MaskSub to CLIP, vision and language, pre-training process. MaskSub is effective for CLIP pre-training.